Given the data:

 

6

8

K

2

4

K

10

3K

 

 The mean of the above data is 9.375 where K is a constant. Find

a)    The value of K.

b)    The coefficient of skewness and interpret.


1
Expert's answer
2021-09-01T19:47:42-0400

"E(X)=( 6+8+K+2+4+K+10+3K)\/8=(30+5K)\/8"

"K=(8E(X)-30)/5=(8\cdot 9.375-30)/5=9"

"E(X^2)=(6^2+8^2+9^2+2^2+4^2+9^2+10^2+27^2)\/8=138.875"

"Var(X)=E(X^2)-E(X)^2=138.875-9.375^2=50.98"

"\sigma(X)=\sqrt{Var(X)}=\sqrt{50.98}=7.14"

Now we calculate Z-scores of each value of the random variable X:

"\tilde X=(X-E(X))/\sigma(X)"

"(6-9.375)\/7.14=-0.4727"

"(8-9.375)\/7.14=-0.1926"

"(9-9.375)\/7.14=-0.0525"

"(2-9.375)\/7.14=-1.0329"

"(4-9.375)\/7.14=-0.7528"

"(9-9.375)\/7.14=-0.0525"

"(10-9.375)\/7.14=0.0875"

"(27-9.375)\/7.14=2.4684"

The coefficient of skewness is by definition

"E(\tilde X^3)=(-0.4727^3-0.1926^3-0.0525^3-1.0329^3-0.7528^3-0.0525^3+0.0875^3+2.4684^3)/8=13.4"

One can see that "E(\tilde X^3)>>0", this means that some values of X (X=27) are located far to the right of the average value. The value "X=27" one should consider as extremal (unusual) big.